OpenStudy (anonymous):

If g(x)=x^2+4x-5, then find g(-4)

jimthompson5910 (jim_thompson5910):

g(x)=x^2+4x-5 g(-4)=(-4)^2+4(-4)-5 g(-4)=16+4(-4)-5 g(-4)=16-16-5 g(-4)=-5